The average kinetic energy of the molecules in a gas sample depends only on the temperature, T. But given the same kinetic energies, a lighter molecule will move faster than a heavier molecule.where R=8.314 J/(mol·K) and ℳ is molar mass in kilograms per mole. Note that a joule is the same as a kg·m2/s2. What is the rms speed of Cl2 molecules at 443 K? What is the rms speed of He atoms at 443 K?
Answer #1

What is the rms speed of Cl2 molecules at 443 K

molar mass of Cl2 = 71 g / mole = 0.071 kg / mole

rms speed of Cl2 = srt (3RT / M) = sqrt (3 * 8.314 * 443 / 0.071) = 394.5 m / s

What is the rms speed of He atoms at 443 K?

molar mass of He = 4.0 g / mole = 0.004 kg / mole

rms speed of He = srt (3RT / M) = sqrt (3 * 8.314 * 443 / 0.004) = 1662 m / s

as He is lighter element in compare to Cl2 and hence  rms speed of He is higher in compare to Cl2.